OpenStudy (anonymous):

There is a charge Q=1 C located at and a charge -Q/2 located at m. What is the E-field (magnitude in units N/C and direction) at m ?

OpenStudy (anonymous):

thnx nurali could u jst help me wid hw1

OpenStudy (anonymous):

A point charge Q_1 = -2 \mbox{ } \mu \mbox{C} is located at x=0, and a point charge Q_2 = +8 \mbox{ } \mu \mbox{C} is placed at x = -0.5 \mbox{ m} on the x-axis of a cartesian coordinate system.The goal of this problem is to determine the electric field, \vec{E}(x) = E(x) \hat{x}, at various points along the x-axis. (a)What is E(x) (in N/C) for x=-64.5 m ? unanswered What is E(x) (in N/C) at x=-0.26 m? unanswered What is E(x) (in N/C) at x=42 m? unanswered (b) At what point (apart from |x| = \infty), is E(x) = 0? Express your answer in meters. unanswered

OpenStudy (anonymous):

first one is wrong

OpenStudy (anonymous):

ya now is correct thnx

OpenStudy (nurali):

Pleasure

OpenStudy (anonymous):

A flat thin non-conducting ring has an inner radius and an outer radius . The disk is uniformly charged with charge per unit area . (a) What is the electric field at the center, , of the ring? Express your answer in terms of the following variables, if necessary, , , and the constant (if needed, enter pi for , epsilon_0 for , sigma for , R_1 for .. etc). unanswered (b)What is the magnitude of the electric field at point which is a distance x above the point ? The line is perpendicular to the plane of the ring. Express your answer in terms of the following variables, if necessary, , , , and the constant (if needed, enter pi for , epsilon_0 for , sigma for , R_1 for .. etc). unanswered

OpenStudy (anonymous):

There are two conducting concentric hollow spheres of outer radii and ( ). The thickness of the material of both spheres is . The inner sphere is negatively charged with charge density . The larger sphere is positively charged with charge density . (a) What is the electric field (magnitude and direction) inside the inner sphere? Direction: use units vectors in spherical coordinates ( , , ) Magnitude: (Express your answer in terms of the following variables, if necessary , , , , , and the constant . To enter type sigma_1, for type sigma_2, for type R_1, etc..) unanswered What is the electric field (magnitude and direction) at distance from the center of the spheres for the following values of : (b) but larger than Direction: use units vectors in spherical coordinates ( , , ) Magnitude: (Express your answer in terms of the following variables, if necessary , , , , , and the constant . To enter type sigma_1, for type sigma_2, for type R_1, etc..) unanswered (c) but smaller than Direction: use units vectors in spherical coordinates ( , , ) Magnitude: (Express your answer in terms of the following variables, if necessary , , , , , and the constant . To enter type sigma_1, for type sigma_2, for type R_1, etc..) unanswered (d) but larger than Direction: use units vectors in spherical coordinates ( , , ) Magnitude: (Express your answer in terms of the following variables, if necessary , , , , , and the constant . To enter type sigma_1, for type sigma_2, for type R_1, etc..) unanswered

OpenStudy (anonymous):

There is a very large horizontal conducting plate in the plane (consider it infinitely large). Its thickness is . The charge densities on the upper and the lower surfaces are both equal to . The goal of this problem is to find the electric field (magnitude and direction) at different points. (a) Consider a point above the upper surface. What is the direction of the electric field? What is the magnitude of the electric field? Express your answer in terms of, if necessary, , and the constant . unanswered (b) Consider now a point below the lower surface. What is the direction of the electric field? What is the magnitude of the electric field? Express your answer in terms of, if necessary, , and the constant . unanswered (c) Consider now a point below the top surface. What is the magnitude of the electric field? Express your answer in terms of, if necessary, , and the constant . unanswered